A store manger orders T-shirts so that’s 15 out of every 35 are a medium. How many medium T-shirts would you expect to find when there are 126 T-shirts on a rack
1
Expert's answer
2019-11-18T13:54:06-0500

Let "x" be the number of medium T-shirts out of 126 on the rack. Write down the ratio of medium T-shirts to the total number of T-shirts:


"\frac{15}{35}=\frac{x}{126}"

Multiply across the known corners, then divide by the third number:


"x=\frac{15\cdot126}{35}""x=54"

Hence, we expect to find 54 T-shirts on a rack.

Answer: 54 T-shirts